# Local Look-Ahead Guidance via Verifier-in-the-Loop for Automated Theorem Proving

### v1

@@ 3.1 Methodology
Our LeanListener framework is based on employing external proof assistant feedback as guidance signal. To bridge the discrepancy between the seq2seq-based training and beam search-based proof full tree expansion during inference, we bring the per-step proof-tree expansion based exploration to the training regime. Aside from transitioning to the next proof state, applying a tactic in Lean results in valuable information, which can be used to score the efficacy of the tactic at the concerned proof step, and consequently guide the model toward proving theorems. For example, a syntactically valid tactic can be gauged for the number of sub-goals in the theorem it helps resolve.
@@
A straightforward approach to do so is employing Direct Preference Optimization (DPO) (Rafailov et al., 2024), a popular lightweight but effective algorithm used to align LLMs with human preferences, which is often replacing reinforcement learning based solutions with separate reward models like RLHF. The key ingredient in DPO is creating a static preference dataset in the form of pairwise comparisons $D=\{(x_{i},y_{i}^{+},y_{i}^{-})\}_{i}$ such that for each prompt $x_{i}$, the generated output $y_{i}^{+}$ is preferred over $y_{i}^{-}$. DPO directly optimizes the policy $\pi_{\theta}$ (possibly, initialized to a reference policy $\pi_{ref}$) via the following loss based on the static dataset:
$L_{DPO}(\pi_{\theta},\pi_{ref})=-\mathbb{E}_{(x,y^{+},y^{-})\sim D}\left[\log\sigma\left(\beta\log\frac{\pi_{\theta}(y^{+}|x)}{\pi_{ref}(y^{+}|x)}-\beta\log\frac{\pi_{\theta}(y^{-}|x)}{\pi_{ref}(y^{-}|x)}\right)\right]$ (1)
where $\sigma$ is the logistic function. The DPO preference dataset can be built based on applicable/not applicable tactics as the positive and negative samples in a pair. In essence, the model can be trained to prefer the applicable tactics, i.e., tactics that lead to a valid proof state, over inapplicable tactics, i.e., tactics that lead to an error state. The unsophisticated DPO pairing configuration will intrinsically reward all the applicable tactics equally, irrespective of whether it takes the proof toward a conclusive state or not. However, a more intricate scoring of tactics can be performed by looking deeper into the Lean feedback, such as the number $\mathcal{G}(s_{t+1})$ of theorem sub-goals that remains to be proven at step $s_{t+1}$. To craft a more differential scoring of applicable tactics, as reward $r_{t}$ for a tactic $a_{t}$, we choose $R(a_{t};s_{t})=\textit{softplus}\left(\mathcal{G}(s_{t})-\mathcal{G}(s_{t+1})\right)$, where $\textit{softplus}(x,\beta)=\frac{1}{\beta}\ln(1+e^{\beta x})$. We set $\beta$ to $0.5$. We invariably score the invalid tactics with a score of 00. Such a scoring strategy rewards, and hence, trains the model to prioritize generating tactics that help resolve more sub-goals.
To take the most out of the Lean feedback and incorporate the number of proven (sub-) goals by the tactics as well, we need a more sophisticated objective. In this regard, we propose to use Group Relative Policy Optimization (GRPO) (Shao et al., 2024), a variant of reinforcement learning (RL) Proximal Policy Optimization (PPO) algorithm (Schulman et al., 2017), in which each tactic is rewarded based on the sub-goals criterion discussed above. For each prompt $q$ (current state $s_{t}$), GRPO samples a group of outputs $\{o_{1},o_{2},\cdots,o_{G}\}$ (generated tactics) from the old policy $\pi_{\theta}{}_{\text{old}}$ and maximizes the policy model with the following objective:
@@
$\min\bigg{[}\frac{\pi_{\theta}(o_{i,t}|q,o_{i,<t})}{\pi_{\theta}{}_{\text{old}}(o_{i,t}|q,o_{i,<t})}\hat{A}_{i,t},\;\text{clip}\bigg{(}\frac{\pi_{\theta}(o_{i,t}|q,o_{i,<t})}{\pi_{\theta}^{\text{old}}(o_{i,t}|q,o_{i,<t})},1-\epsilon,1+\epsilon\bigg{)}\hat{A}_{i,t}\bigg{]}$
@@
$D_{\text{KL}}$ $(\pi_{\theta}\,||\,\pi_{\text{ref}})=\frac{\pi_{\text{ref}}(o_{i,t}|q,o_{i,<t})}{\pi_{\theta}(o_{i,t}|q,o_{i,<t})}-\log\left(\frac{\pi_{\text{ref}}(o_{i,t}|q,o_{i,<t})}{\pi_{\theta}(o_{i,t}|q,o_{i,<t})}\right)-1$
We use the pre-trained ReProver model from Yang et al. (2023) both as a base reference model for $\pi_{ref}$ and to initialize the policy to optimize $\pi_{\theta}$.
3.2 Dataset Curation
For both DPO and GRPO finetuning, we build our training dataset using the training split of LeanDojo Benchmark (Yang et al., 2023) to avoid data contamination as the ReProver model has already been pre-trained on it. Inspired by the online-RL training paradigms for eliciting reasoning in LLMs (Gloeckle et al., 2024), we, too, build our training paradigm in an online setting. During online training, we update the model used in the beam search to generate tactic proposals for every $50$ training iteration with the target model. We also experiment with the offline setting, where we use the pre-trained ReProver model to generate the training data offline. We discuss the relative advantages of online over offline training in Table 1 and Sec. 4.
To generate the fine-tuning data, we parse the human-annotated proofs in the LeanDojo benchmark using either the reference ReProver model (offline) or the target model (online). More specifically, for each proof state $s_{t}$ in human-annotated theorem proof in training split, we sample $w=8$ tactics $\{a_{t}^{j}\sim\pi_{ref}(a|s_{t})\}_{j}^{w}$ via beam search and sort them by their likelihood score given by the sampling model. Additionally, we look for the presence of the ground truth tactic, i.e., gold tactic, in the beam search proposals: if the golden tactic $\widehat{a_{t}}$ for the current state $s_{t}$ happens to be absent from the top $w$ proposals, we append it as a $w+1=9$-th tactic at the end of the list (i.e. with lowest likelihood); this means states can have either $8$ or $9$ tactic proposals for training. The next step involves accruing feedback from the Lean by individual application of each proposed tactic on the current proof state.
@@ 3.2 Dataset Curation
Table 1: This table reports different metrics to measure the step-wise performance. Prec.@$8$: percentage of valid tactics among the top $8$ tactics sampled at each proof state. MAP: mean average precision percentage. MRR: percentage of mean reciprocal rank. Len. Valid Tactics: average length (in terms of tokens) of the valid generated tactics. Len. All Tactics: average length of all generated tactics. %00-precision steps: fraction of states with no valid tactics.
@@
For the GRPO scenario, the dataset is being built dynamically throughout the training. More specifically, for each state in the training set, the model generates a set of tactics using beam search. Then, using the reward function described in Sec. 3.1, each generated tactic is scored. The more (sub)goals a tactic solves, the greater its reward. Since the policy model is being updated through GRPO objective, the model’s generated tactics change over time. Similar to the online case in DPO, we update the model used for sampling tactics every $50$ training steps with the on-policy target model.
3.3 Training Setting
@@ 4 Experiments
We study the effectiveness of LeanListener in different aspects on the test set of LeanDojo benchmark. The test set has two variants, random, where the whole dataset is randomly split into training, validation, and test set, and the more challenging novel premises, where the test set includes premises that do not appear in the training222 Yang et al. (2023) only released the model trained on the random split, which we use as base for our method. We report the performance of the methods also on the novel premises test despite the possible overlap with the training set since this testset is still more challenging than the random one. See github issue.. In the following experiments, we first verify whether incorporating Lean feedback enhances the quality of generated tactics. Then, we examine the performance on theorem proving. Finally, we study the improved theorem proving speed of our method.
@@
To begin with, we investigate to what extent adding DPO and GRPO during the training pipeline increases the number of valid tactics generated by the model. A tactic is valid if it is applicable to a state by Lean, even if it does not lead to a proof. The validity is also about the syntactic correctness of the tactic. We expect that if the model generates more valid tactics in the best first search, the chance of proving the given theorem, as well as the efficiency of the proof, will increase.
Table 1 presents the numerical results on the validity of the generated tactics under different scenarios using DPO and GRPO. In this table, we compare the state-wise metrics which we expect to correlate with our training objective. First of all, using the DPO objective, we observe that all offline strategies fail to improve the number of valid tactics generated by the model, while the online strategy significantly improves the number of valid tactics, and reduces the number of states with no valid tactics. Online GRPO provides the best results. It is interesting to observe that the length of valid tactics decrease in the best models, thereby, reducing unnecessarily complex tactic generation.
@@
Moving forward, we investigate whether improved tactic generation results in stronger reasoning capabilities or not. Table 2 presents the final theorem proving the performance of our models with the pre-trained ReProver baseline along with GPT-4 (Achiam et al., 2023) and tidy ( which is a non-machine learning and heuristic-based approach), on the LeanDojo benchmark when using the random split. We first observe that while DPO training on binary feedback improves the number of valid tactics, it always leads to a small degradation in the number of proven theorems. We believe that there is spurious correlation between the length of the tactics and their validity on the paired samples in DPO training, which leads to a bias toward generating longer and not necessarily useful tactics. GRPO addresses effectively this issue by online training and avoiding paired samples for training, as well as employing more informative feedback about the number of solved sub-goals. Indeed, we find that this finer reward, which incorporates the number of solved sub-goals, is more effective and leads to an increased number of proven theorems.
@@
In the previous experiments, we show that LeanListener generates a more efficient search tree by having more valid tactics. Now, we want to evaluate if this approach leads to a more efficient inference time. Figure LABEL:fig:inference_time compares the inference time in LeanListener and ReProver as the baseline model. As can be observed, LeanListener proves more theorems in up to $20\%$ less time.
@@
In addition to the aggregate Pass@1 performance metric discussed earlier, we present an analysis of the proof lengths generated by the LeanListener and ReProver models. Generally, concise proofs, i.e., those with fewer proof steps, are considered more efficient and desirable than lengthier ones. Our aim here is to provide insights into the quality of proofs generated by both models. As shown in Figure LABEL:fig:proof_len, our LeanListener consistently produces more concise proofs compared to the base ReProver model.
@@ 5 Conclusions
In this paper, we propose a novel RL-based framework, LeanListener, based on the provided feedback by an external automatic verifier to improve trained LLMs in automatic theorem proving. LeanListener considers the general applicability of tactics as well as their local effectiveness, i.e., their impact on the number of unproven (sub)goals in a proof sequence, to fine-tune theorem provers. Such step-wise reinforcement with the corresponding reward efficiently harnesses the benefits of the trajectory level guidance without outcome-dependent reward computation. Our experimental results show that LeanListener not only surpasses the considered baselines in proving more theorems but also does so faster in less inference time.
@@ Appendix A Extended Discussions on Related Works
GPT-f [Polu and Sutskever, 2020] Metamath Prover Proof Step - Best First Search
Lean-Gym [Polu et al., 2023] Lean Prover Proof Step - Best First Search Expert iteration
LeanDojo [Yang et al., 2023] Lean Prover Proof step RAG Offline Database Best First Search
LeanAgent [Kumarappan et al., 2024] Lean Prover Proof step RAG Online Database Best First Search
HTPS [Lample et al., 2022] Lean Metamath Prover Proof step - HTPS
DeepSeek-Prover [Xin et al., 2024a] Lean Prover Reasoning Whole Proof - -
DeepSeek-Prover 1.5 [Xin et al., 2024b] Lean Prover Reasoning Proof Step Whole Proof - MCTS
LLEMA [Azerbayev et al., 2024] Any Tool Prover Reasoning Proof Text - Best First Search
Verify Step by Step [Lightman et al., 2023] ORM-PRM Reasoning Text - Best First Search
Table 3: A summary of models for theorem proving and mathematical reasoning.
@@
Machine learning is currently used in automatic theorem proving to generate the whole proof, or provide assistance in sequential proofs by finding the right premise or right tactics, or assisting in theorem prover specific formalization of mathematical statements, as well as combination of these ideas; see [Xin et al., 2024a, b, Polu and Sutskever, 2020, Jiang et al., 2022a, Wang et al., 2023, Jiang et al., 2022b, Wu et al., 2022] for some pointers. We review some of these works in more details. A key component for theorem proving is the formal proof management systems and verifiers like Lean [Moura and Ullrich, 2021] and Coq [Coq Development Team, 2024]. Lean has received particular attention in the mathematics community for example by its use in verification of some components in the condensed mathematics research program [Scholze, 2022]. We use Lean in this work, and a framework that consists of a premise retrieval part and a prover that generates step-wise proof tactics.
Machine learning can be used to find useful premises, definitions and theorems [Irving et al., 2016]. For Coq framework [Coq Development Team, 2024], the authors in Blaauwbroek et al. [2024] introduce a graph neural network for embedding the new definitions using a hierarchical representation based on graph representations of theorems and definitions in the Tactician platform [Blaauwbroek, 2024]. This enables them to use recent proofs and theorems in Coq, while kNN is used for the more recent tactics written by the user. In Thor [Jiang et al., 2022a], the authors introduce a class of methods to use automated theorem provers for premise selection. LeanDojo [Yang et al., 2023] provided a framework that retrieves the premise from a database of Lean premises and uses ReProver to generate tactics for theorem proving in Lean. LeanAgent [Kumarappan et al., 2024] adds a dynamic database to LeanDojo, which enables the continual learning of the agent. Besides, a new curriculum learning based on the difficulty of the theorems is also added.
@@
Searching over different proofs and tactics is another component explored in various works, for example [Polu et al., 2023, Xin et al., 2024b, Lample et al., 2022]. As discussed in Polu et al. [2023], any RL algorithm for theorem proving should address two challenges, namely an infinite dimensional action space, and the absence of a natural opponent for self-play. Therefore, the authors suggest using expert iteration [Anthony et al., 2017], which amounts to iteratively fine-tuning a based model and searching to generate correct proofs. They also introduce lean-gym to facilitate the search procedure in Lean. HTPSs was introduced in Lample et al. [2022] which is a new search algorithms within an online reinforcement learning procedure. The key idea is that the search is represented over a hypergraph, where a policy network generates tactics composed of sub-goals, and then each goal is expanded for proof using new tactics. The provability of each goal is approximated using a critic. The idea of keeping the visit counts, action values and their statistics follows the similar MCTS procedure. In Gloeckle et al. [2024], the authors introduce a reinforcement learning based framework for theorem proving in Lean 4, which consist of using a programming interface based on Aesop for proof search organization, the HTPS ([Lample et al., 2022]) procedure with an online reinforcement learning step. In Zhao et al. [2023], the authors use subgoal learning from reinforcement learning to decompose an informal LLM generated proof into subgoals and verify using a verifier. They also use a diffusion model for demonstration organization.
@@
The transformer based language models are used also for theorem proving. For example GPT-f was introduced in Polu and Sutskever [2020] using Metamath as the formalization framework. Besides, they showed that iterative training a value functionon proofs generated by the prover can continually improve the performance. Another example is LLEMMA which is based on pretrained Llama Code [Azerbayev et al., 2024].
@@
The proofs in some frameworks are already in the language of formal theorem provers. Besides, the external tools for mathematical reasoning can include codes for mathematical arguments. Machine learning has been used to help the formalization of mathematical statements suitable for the automatic theorem provers [Wu et al., 2022, Jiang et al., 2022b, Zhao et al., 2023]. Furthermore, there is a spectrum of methods that starts here with methods operating in a formal language and ends with informal proofs and solutions in natural language. In MathCoder2 Lu et al. [2024a], the authors provided pairs of mathematical codes and the associated natural language versions.
@@
DeepSeek-Prover [Xin et al., 2024a] generated Lean4 proof data by translating natural language problems into their formal versions in Lean4. The model produces the whole proof in a single turn. DeepSeekV1.5 [Xin et al., 2024b] provides a middle ground by generating the whole proof, verifying it via Lean, and then truncating the proof until the first error. The authors propose additional techniques relying on proper appending of previous states and including truncate-and-resume in the MCTS procedure. They also leverage Verification feedback from Lean and improve model’s alignment with the formal structure of Lean. The authors in Wang et al. [2023] provide another hybrid solution where first an informal proof is generated, then broken into sub-components, and then a lemma is retrieved from a library of lemmas considered as skills, and finally the final formalized proof is generated using the previous components. They use the framework of Isabelle [Paulson, 1994] for theorem proving.
@@
Treating mathematical problem solving as a reasoning task has been considered in many works, for example [Tong et al., 2024, Zhu et al., 2023, Shao et al., 2022, 2024, Guan et al., 2025]. Among them, there are models, like Shao et al. [2024], based on chain-of-thought [Wei et al., 2022] or program of thought [Gao et al., 2023, Chen et al., 2022]. The generator-verifier configuration has been considered in Zhu et al. [2023], where they train a step and path verifier for reasoning. Integrating a tool for improving reasoning of large language models has been used in Gou et al. [2023]. In contrast to our work, these models do not rely on theorem provers. The authors in Wang and Deng [2020] address the lack of human labeled theorems and proofs to use for supervised training by training a generative model and, then, use the synthesized pairs to train a theorem prover model.
@@
The idea of proof size and using it for guiding the search has been discussed in Wu et al. [2021], Polu et al. [2023]. There is a line of work related to discovering new mathematical functions or solving mathematical problems, see for example [Real et al., 2023, Alfarano et al., 2024, Marchetti et al., 2023, Wenger et al., 2022]. However, these works do not rely on proof verifiers and fall outside the scope of this work.

### v2

@@ 3.2 LeanListener
Our LeanListener framework is based on employing external proof assistant feedback as a guidance signal. To bridge the discrepancy between the seq2seq-based training and beam search-based proof full tree expansion during inference, we bring the per-step proof-tree expansion based exploration to the training regime. Aside from transitioning to the next proof state, applying a tactic in Lean results in valuable information, which can be used to score the efficacy of the tactic at the concerned proof step, and consequently guide the model toward proving theorems. For example, a syntactically valid tactic can be gauged for the number of sub-goals in the theorem it helps resolve.
@@
A straightforward approach to do so is employing Direct Preference Optimization (DPO) Rafailov et al. (2024), a popular lightweight but effective algorithm used to align LLMs with human preferences, which is often replacing reinforcement learning based solutions with separate reward models like RLHF. The key ingredient in DPO is creating a static preference dataset in the form of pairwise comparisons $D=\{(x_{i},y_{i}^{+},y_{i}^{-})\}_{i}$ such that for each prompt $x_{i}$, the generated output $y_{i}^{+}$ is preferred over $y_{i}^{-}$. DPO directly optimizes the policy $\pi_{\theta}$ (possibly, initialized to a reference policy $\pi_{ref}$) via the following loss based on the static dataset:
$L_{DPO}(\pi_{\theta},\pi_{ref})=-\mathbb{E}_{(x,y^{+},y^{-})\sim D}$ (1)
$\left[\log\sigma\left(\beta\log\frac{\pi_{\theta}(y^{+}|x)}{\pi_{ref}(y^{+}|x)}-\beta\log\frac{\pi_{\theta}(y^{-}|x)}{\pi_{ref}(y^{-}|x)}\right)\right]$
where $\sigma$ is the logistic function. The DPO preference dataset can be built based on applicable/not applicable tactics as the positive and negative samples in a pair. In essence, the model can be trained to prefer the applicable tactics, i.e., tactics that lead to a valid proof state, over inapplicable tactics, i.e., tactics that lead to an error state. The unsophisticated DPO pairing configuration will intrinsically reward all the applicable tactics equally, irrespective of whether it takes the proof toward a conclusive state or not. However, a more intricate scoring of tactics can be performed by looking deeper into the Lean feedback, such as the number $\mathcal{G}(s_{t+1})$ of theorem sub-goals that remains to be proven at step $s_{t+1}$. To craft a more differential scoring of applicable tactics, as reward $r_{t}$ for a tactic $a_{t}$, we choose $R(a_{t};s_{t})=\textit{softplus}\left(\mathcal{G}(s_{t})-\mathcal{G}(s_{t+1})\right)$, where $\textit{softplus}(x,\beta)=\frac{1}{\beta}\ln(1+e^{\beta x})$. We invariably score the invalid tactics with a score of 00. Such a scoring strategy rewards, and hence, trains the model to prioritize generating tactics that help resolve more sub-goals. We choose softplus over a naive linear delta to prevent negative reward values, since the number of intermediate sub-goals may temporarily increase before the proof is completed.
To take the most out of the Lean feedback and incorporate the number of proven (sub-) goals by the tactics as well, we need a more sophisticated objective. In this regard, we propose to use Group Relative Policy Optimization (GRPO) Shao et al. (2024), a variant of reinforcement learning (RL) Proximal Policy Optimization (PPO) algorithm Schulman et al. (2017), in which each tactic is rewarded based on the sub-goals criterion discussed above. For each prompt $q$ (current state $s_{t}$), GRPO samples a group of outputs $\{o_{1},o_{2},\cdots,o_{G}\}$ (generated tactics) from the old policy $\pi_{\theta}{}_{\text{old}}$ and maximizes the policy model with the following objective:
@@
$\min\bigg{[}\frac{\pi_{\theta}(o_{i,t}|q,o_{i,<t})}{\pi_{\theta}{}_{\text{old}}(o_{i,t}|q,o_{i,<t})}\hat{A}_{i,t},$
$\;\text{clip}\bigg{(}\frac{\pi_{\theta}(o_{i,t}|q,o_{i,<t})}{\pi_{\theta}^{\text{old}}(o_{i,t}|q,o_{i,<t})},1-\epsilon,1+\epsilon\bigg{)}\hat{A}_{i,t}\bigg{]}$
@@
$D_{\text{KL}}$ $(\pi_{\theta}\,||\,\pi_{\text{ref}})=$ (3)
$\frac{\pi_{\text{ref}}(o_{i,t}|q,o_{i,<t})}{\pi_{\theta}(o_{i,t}|q,o_{i,<t})}-\log\left(\frac{\pi_{\text{ref}}(o_{i,t}|q,o_{i,<t})}{\pi_{\theta}(o_{i,t}|q,o_{i,<t})}\right)-1$
We use the pre-trained generator in the ReProver model from Yang et al. (2023) both as a base reference model for $\pi_{ref}$ and to initialize the policy to optimize $\pi_{\theta}$.
@@
Table 1: The DPO objective optimizes the model to prefer valid tactics. In this table, we report different metrics to measure the step-wise performance. Prec.@$8$ = the percentage of the number of valid tactics among the top $8$ tactics sampled at each proof state. MAP= percentage of mean average precision. MRR = percentage of mean reciprocal rank. Len. Valid Tactics = average length (in terms of tokens) of the generated tactics which are valid. Len. All Tactics = average length of all generated tactics. %00-precision steps = fraction of states with no valid tactics.
3.2.1 Dataset Curation
For both DPO and GRPO finetuning, we build our training dataset using the training split of LeanDojo Benchmark Yang et al. (2023) to avoid data contamination as the ReProver model has already been pre-trained on it. Inspired by the online-RL training paradigms for eliciting reasoning in LLMs Gloeckle et al. (2024), we, too, build our training paradigm in an online setting. During online training, we update the model used in the beam search to generate tactic proposals for every $50$ training iteration with the target model. We also experiment with the offline setting, where we use the pre-trained ReProver model to generate the training data offline. We discuss the relative advantages of online over offline training in Table 1 and Sec. 4.
To generate the fine-tuning data, we parse the human-annotated proofs in the LeanDojo benchmark using either the reference ReProver model (offline) or the target model (online). More specifically, for each proof state $s_{t}$ in human-annotated theorem proof in training split, we sample $w=8$ tactics $\{a_{t}^{j}\sim\pi_{ref}(a|s_{t})\}_{j}^{w}$ via beam search and sort them by their likelihood score given by the sampling model. Additionally, we look for the presence of the ground truth tactic, i.e., gold tactic, in the beam search proposals: if the golden tactic $\widehat{a_{t}}$ for the current state $s_{t}$ happens to be absent from the top $w$ proposals, we append it as a $w+1=9$-th tactic at the end of the list (i.e. with lowest likelihood); this means states can have either $8$ or $9$ tactic proposals for training. The next step involves accruing feedback from the Lean framework by individual application of each proposed tactic on the current proof state.
@@ 3.2.1 Dataset Curation
For the GRPO scenario, the dataset is being built dynamically throughout the training. More specifically, for each state in the training set, the model generates a set of tactics using beam search. Then, using the reward function described in Sec. 3.2, each generated tactic is scored. The more (sub)goals a tactic solves, the greater its reward. Since the policy model is getting updated through GRPO objective, the model’s generated tactics change over time. Similar, to the online case in DPO, we update the model used for sampling tactics every $50$ training iterations with the on-policy target model.
3.2.2 Training Setting
@@
Figure 3: To understand the improvement in inference speed, in these figures, we visualize the number of theorems in the two test sets which are proven by each model within different inference time limits. While the overall number of theorems proved within $10$ minutes by the two models is comparable, we see LeanListener is consistently faster, which means, it takes significantly less time to solve the same number of theorems (see the purple lines for example).
@@ 4 Experiments
We evaluate the effectiveness of LeanListener in different aspects on the test set of LeanDojo benchmarks. The test set has two variants, random, where the whole dataset is randomly split into training, validation, and test set, and the more challenging one, novel premises, where the test set includes premises that do not appear in the training data222 Yang et al. (2023) only released the model trained on the random split, which we use as the base for our method. We report the performance of the methods also on the novel premises test despite the possible overlap with the training set since this testset is still more challenging than the random one. See github issue.. In the following experiments, we first verify whether incorporating Lean feedback enhances the quality of generated tactics. Then, we examine the performance on theorem proving task. Finally, we study the improved theorem proving speed of our method.
@@
To begin with, we investigate to what extent adding DPO and GRPO in the training pipeline increases the number of valid tactics generated by the model. As mentioned before, a tactic is valid if it is applicable to a state by Lean, even if it does not lead to a proof. The validity is also about the syntactic correctness of the tactic. We expect that if the model generates more valid tactics in the best first search, the chance of proving the given theorem, as well as the efficiency of the proof, will increase.
Table 1 presents the numerical results on the validity of the generated tactics under different scenarios using DPO and GRPO objectives. In this table, we compare the state-wise metrics which we expect to correlate with our training objective. First of all, using the DPO objective, we observe that all offline strategies fail to improve the number of valid tactics generated by the model, while the online strategy significantly improves the number of valid tactics, and reduces the number of states with no valid tactics. Among all, online GRPO provides the best results. It is interesting to observe that the length of valid tactics decreases in the best models, thereby reducing unnecessarily complex tactic generation.
@@
Moving forward, we investigate whether improved tactic generation results in stronger reasoning capabilities or not. Table 2 presents the final theorem proving the performance of our models with the pre-trained ReProver baseline along with GPT-4 Achiam et al. (2023) and tidy (which is a non-machine learning and heuristic-based approach), on the LeanDojo benchmark when using the random split.  [^3] The evaluation results on the MiniF2F Zheng et al. (2022) and ProofNet Azerbayev et al. (2023) datasets are reported in Appendix C.. We first observe that while DPO training on binary feedback improves the number of valid tactics, it always leads to a small degradation in the number of proven theorems. We believe that there is a spurious correlation between the low-ranked valid tactics and their length in the paired samples in DPO training, which leads to a bias toward generating longer and not necessarily useful tactics. GRPO effectively addresses this issue by online training and avoiding paired samples for training, as well as employing more informative feedback about the number of solved sub-goals. Indeed, we find that this finer reward, which incorporates the number of solved sub-goals, is more effective and leads to an increased number of proven theorems.
@@
In the previous experiments, we show that LeanListener generates a more efficient search tree by having more valid tactics. Now, we want to evaluate if this approach leads to a more efficient inference time. Figure 3 compares the inference time in LeanListener and ReProver as the baseline model on the LeanDojo benchmark. As can be observed, LeanListener proves more theorems in up to $20\%$ less time for both random and novel premises splits.
@@
In addition to the aggregate Pass@1 performance metric discussed earlier, we present an analysis of the proof lengths generated by the LeanListener and ReProver models on the LeanDojo benchmark [^4] See Appendix C for the analysis on MiniF2F and ProofNet.. Generally, concise proofs, i.e., those with fewer proof steps, are considered more efficient and desirable than lengthier ones. Our aim here is to provide insights into the quality of proofs generated by both models. As shown in Figure LABEL:fig:proof_len, our LeanListener consistently produces more concise proofs compared to the base ReProver model.
@@ 5 Conclusions
In this paper, we propose a novel RL-based framework, LeanListener, based on the provided feedback by an external automatic verifier to improve trained LLMs in automatic theorem proving. LeanListener considers the general applicability of tactics as well as their local effectiveness, i.e., their impact on the number of unproven (sub)goals in a proof sequence, to fine-tune theorem provers. Such step-wise reinforcement with the corresponding reward can efficiently harness the benefits of the trajectory level guidance without outcome dependent reward computation. Our experimental results show that LeanListener not only surpasses the considered baselines in proving more theorems on the LeanDojo benchmark but also does so faster in less inference time.
6 Limitations
In this work, we have focused on a well-known open-source ITP model to show the effectiveness of the LeanListener framework. Applying our framework to more tactic generator models with different architecture, model sizes, and components in the training pipelines is an interesting future direction. We also mainly study the performance of LeanListener on the LeanDojo benchmark. Investigating the performance of LeanListener on other theorem proving datasets is another potential follow-up work.
@@ Appendix A Extended Discussions on Related Works
GPT-f Polu and Sutskever (2020) Metamath Prover Proof Step - Best First Search
Lean-Gym Polu et al. (2023) Lean Prover Proof Step - Best First Search Expert iteration
LeanDojo Yang et al. (2023) Lean Prover Proof step RAG Offline Database Best First Search
LeanAgent Kumarappan et al. (2024) Lean Prover Proof step RAG Online Database Best First Search
HyperTree Proof Search (HTPS) Lample et al. (2022) Lean Metamath Prover Proof step - HTPS
DeepSeek-Prover Xin et al. (2024a) Lean Prover Reasoning Whole Proof - -
DeepSeek-Prover 1.5 Xin et al. (2024b) Lean Prover Reasoning Proof Step Whole Proof - MCTS
LLEMA Azerbayev et al. (2024) Any Tool Prover Reasoning Proof Text - Best First Search
Verify Step by Step Lightman et al. (2023) ORM-PRM Reasoning Text - Best First Search
Table 3: A summary of models for theorem proving and mathematical reasoning
@@
Machine learning is currently used in automatic theorem proving to generate the whole proof, or provide assistance in sequential proofs by finding the right premise or right tactics, or assisting in theorem prover specific formalization of mathematical statements, as well as combination of these ideas; see Xin et al. (2024a, b); Polu and Sutskever (2020); Jiang et al. (2022a); Wang et al. (2023); Jiang et al. (2022b); Wu et al. (2022) for some pointers. We review some of these works in more details. A key component for theorem proving is the formal proof management systems and verifiers like Lean Moura and Ullrich (2021) and Coq Coq Development Team (2024). Lean has received particular attention in the mathematics community for example by its use in verification of some components in the condensed mathematics research program Scholze (2022). We use Lean in this work, and a framework that consists of a premise retrieval part and a prover that generates step-wise proof tactics.
@@
Searching over different proofs and tactics is another component explored in various works, for example Polu et al. (2023); Xin et al. (2024b); Lample et al. (2022). As discussed in Polu et al. (2023), any RL algorithm for theorem proving should address two challenges, namely an infinite dimensional action space, and the absence of a natural opponent for self-play. Therefore, the authors suggest using expert iteration Anthony et al. (2017), which amounts to iteratively fine-tuning a based model and searching to generate correct proofs. They also introduce lean-gym to facilitate the search procedure in Lean. HTPSs was introduced in Lample et al. (2022) which is a new search algorithms within an online reinforcement learning procedure. The key idea is that the search is represented over a hypergraph, where a policy network generates tactics composed of sub-goals, and then each goal is expanded for proof using new tactics. The provability of each goal is approximated using a critic. The idea of keeping the visit counts, action values and their statistics follows the similar MCTS procedure. In Gloeckle et al. (2024), the authors introduce a reinforcement learning based framework for theorem proving in Lean 4, which consist of using a programming interface based on Aesop for proof search organization, the HTPS (Lample et al. (2022)) procedure with an online reinforcement learning step. In Zhao et al. (2023), the authors use subgoal learning from reinforcement learning to decompose an informal LLM generated proof into subgoals and verify using a verifier. They also use a diffusion model for demonstration organization.
@@
The transformer based language models are used also for theorem proving. For example GPT-f was introduced in Polu and Sutskever (2020) using Metamath as the formalization framework. Besides, they showed that iterative training a value functionon proofs generated by the prover can continually improve the performance. Another example is LLEMMA which is based on pretrained Llama Code Azerbayev et al. (2024).
@@
The proofs in some frameworks are already in the language of formal theorem provers. Besides, the external tools for mathematical reasoning can include codes for mathematical arguments. Machine learning has been used to help the formalization of mathematical statements suitable for the automatic theorem provers Wu et al. (2022); Jiang et al. (2022b); Zhao et al. (2023). Furthermore, there is a spectrum of methods that starts here with methods operating in a formal language and ends with informal proofs and solutions in natural language. In MathCoder2 Lu et al. (2024a), the authors provided pairs of mathematical codes and the associated natural language versions.
@@
DeepSeek-Prover Xin et al. (2024a) generated Lean4 proof data by translating natural language problems into their formal versions in Lean4. The model produces the whole proof in a single turn. DeepSeekV1.5 Xin et al. (2024b) provides a middle ground by generating the whole proof, verifying it via Lean, and then truncating the proof until the first error. The authors propose additional techniques relying on proper appending of previous states and including truncate-and-resume in the MCTS procedure. They also leverage Verification feedback from Lean and improve model’s alignment with the formal structure of Lean. The authors in Wang et al. (2023) provide another hybrid solution where first an informal proof is generated, then broken into sub-components, and then a lemma is retrieved from a library of lemmas considered as skills, and finally the final formalized proof is generated using the previous components. They use the framework of Isabelle Paulson (1994) for theorem proving.
@@
Treating mathematical problem solving as a reasoning task has been considered in many works, for example Tong et al. (2024); Zhu et al. (2023); Shao et al. (2022, 2024); Guan et al. (2025). Among them, there are models, like Shao et al. (2024), based on chain-of-thought Wei et al. (2022) or program of thought Gao et al. (2023); Chen et al. (2022). The generator-verifier configuration has been considered in Zhu et al. (2023), where they train a step and path verifier for reasoning. Integrating a tool for improving reasoning of large language models has been used in Gou et al. (2023). In contrast to our work, these models do not rely on theorem provers. The authors in Wang and Deng (2020) address the lack of human labeled theorems and proofs to use for supervised training by training a generative model and, then, use the synthesized pairs to train a theorem prover model.
@@
The idea of proof size and using it for guiding the search has been discussed in Wu et al. (2021); Polu et al. (2023). There is a line of work related to discovering new mathematical functions or solving mathematical problems, see for example Real et al. (2023); Alfarano et al. (2024); Marchetti et al. (2023); Wenger et al. (2022). However, these works do not rely on proof verifiers and fall outside the scope of this work.